What is 8x+2 =5x+23 using the balance method?
Answer by Simnepi(216)

to solve any type of equation with an unknown on both sides of the equals sign....
look for the side with the least unknowns
so 8x+2=5x+23
the right hand side (of the equals sign) has 5x and the left hand side has 8x.
so we take 5x from both sides to eliminate the +5x from the right hand side
giving
8x+2-5x=5x+23-5x
simplifying gives
3x+2=23
Now we have to eliminate the number from the side with the unknown on
So we have to subtract 2 from both sides to get rid of the +2 from the right hand side
giving
3x+2-2=23-2
simplifying gives
3x=21
Dividing both sides by 3 gives us
x=7